# Soliton resuscitations: asymmetric revivals of the breathing mode of an atomic bright soliton in a harmonic trap

### v3

@@ III.3.2 Normalization
The less obvious phenomenon that we have seen in this paper, however, is that these expected breathing revivals—our “resuscitations”— have an asymmetric profile, which becomes more pronounced in later resuscitations. The breathing amplitude actually starts to revive distinctly earlier than the first $\pi/\omega$ half-period, and this early revival becomes steadily stronger in later resuscitations. Almost immediately after the half-period, then, the breathing amplitude falls sharply, before more gradually decaying until the next resuscitation.
@@
Qualitatively, we can explain this phenomenon simply, from the spectral facts that we have described above. The two lowest modes of the Bogoliubov-de Gennes spectrum remain as the time-translation zero mode and the collective dipole mode with eigenfrequency exactly $\omega$, while all the other modes are raised above a large gap $\hbar\kappa^{2}/(2M)$, because excitations out of the condensate must all be excited above the large mean-field binding energy of the soliton. When the soliton is much narrower than the trap width, the spectrum above this chemical potential gap is still approximately like that of a single harmonic oscillator. Nevertheless the entire above-gap spectrum has to be slightly distorted away from the even level spacing of an oscillator, to “replace” the missing two lowest modes. These distortions are indeed supplied by the differences between the Bogoliubov-de Gennes equations and the Schrödinger equation of a single trapped particle.

### v4

@@ IV.3.4 Evaluating $B_{N}(\Delta t)$ for $1\ll N\ll\varepsilon^{-2}$
The forms of the envelopes of the breathing amplitude at successive resuscitations are the more surprising phenomenon that we have seen in this paper. These breathing revivals have an asymmetric profile, which becomes more pronounced in later resuscitations. The breathing amplitude actually starts to revive distinctly earlier than the first $\pi/\omega$ half-period, and this early revival becomes steadily stronger in later resuscitations. Almost immediately after the half-period, then, the breathing amplitude falls sharply, before more gradually decaying until the next resuscitation.
@@
Qualitatively, we can explain the asymmetrical resuscitation phenomenon simply, from the spectral facts that we have described above. The two lowest modes of the Bogoliubov-de Gennes spectrum remain as the time-translation zero mode and the collective dipole mode with eigenfrequency exactly $\omega$, while all the other modes are raised above a large gap $\hbar\kappa^{2}/(2M)$, because excitations out of the condensate must all be excited above the large mean-field binding energy of the soliton. When the soliton is much narrower than the trap width, the spectrum above this chemical potential gap is still approximately like that of a single harmonic oscillator. Nevertheless the entire above-gap spectrum has to be slightly distorted away from the even level spacing of an oscillator, to “replace” the missing two lowest modes. These distortions are indeed supplied by the differences between the Bogoliubov-de Gennes equations and the Schrödinger equation of a single trapped particle.
@@
V.5 V.E Analogous effects in higher dimensions?
The small parameter $\varepsilon=1/(\kappa a_{0})$, the ratio of soliton width to trap width, has been important in enabling all our calculations. It is also essential conceptually, however, for our view of the soliton breathing mode as an open quantum system in a non-Markovian environment. If the soliton width and the trap width were comparable, we would simply have a trapped condensate with an even collective mode; there would be no basis for the open-system picture, and the multiple time scales of breathing, attenuation, and revival would not exist.
The length scale hierarchy $\kappa^{-1}\ll a_{0}$, and the associated time scale hierarchy $\omega\ll\hbar\kappa^{2}/(2M)$, exist in this quasi-one-dimensional scenario only because in one dimension a bright soliton defines its own width, through the attractive nonlinearity, independently of any trapping confinement. This kind of self-confinement independently from a trap does not occur in more than one dimension, but one might consider radial breathing of a vortex core as kind of inverted analog of our scenario, in which it is a small void within a much larger condensate that breathes, instead of a small condensate in large empty trap.
Collective excitations of vortex cores in quasi-two-dimensional condensates with attractive interactions have been studied in [19], where in particular a regime of large-amplitude collective excitations is found in which a condensate with a vortex repeatedly breaks up into two clouds that then recombine. While our breathing mode with emission of atoms can be considered a smaller-amplitude version of this behavior, the fact that in the 2D scenario the whole condensate, the vortex core, and the trap width are all of comparable scale prohibits the kind of multi-scale evolution that can be seen as short-period breathing with attenuation and revivals over a much longer period. The prospects for higher-dimensional analogs to our resuscitations seem to be inherently limited for attractively interacting condensates.
In repulsively interacting condensates, on the other hand, a large condensate can have a vortex with a much smaller core, since the core size is set by the interacting strength and the local condensate density. A breathing mode of a vortex core in a repulsively interacting condensate has been studied in [20], with frequencies found to be on the order of the chemical potential scale, as it is in our soliton case. The time-dependent variational Ansatz approach used in [20] assumes that core radius oscillation is a single collective mode, however, and so does not allow for the possibility that it is a component in a quasi-continuum of many modes and therefore decays through dispersion. Hence it is not clear from [20] whether vortex core breathing might attenuate and then revive as our soliton breathing mode does. It may in future be worth performing a similar analysis to the one that we have offered in this paper, based on identifying breathing as a component in a complete set of collective modes, for vortex cores in large, repulsively interacting condensates in two or three dimensions.